OpenStudy (crashonce):

If p=1+2-3-4+5+6-7-8+...+1986, then p equals?

OpenStudy (crashonce):

you must use AP/GP

OpenStudy (crashonce):

@ganeshie8 @zepdrix @iGreen

TheSmartOne (thesmartone):

Did you know that 1+2+3+...10= (1+10)*(10/2)=55

TheSmartOne (thesmartone):

So using that we get (1+1986)*(1986/2)=??

OpenStudy (anonymous):

Sum = (\[Sum = \frac{ n (a + l)}{ 2 }\] Sum = (\[Sum = \frac{ 1986 (1 + 1986)}{ 2 }\] Sum = 1973091

TheSmartOne (thesmartone):

This is because we get pairs